OpenStudy (anonymous):

A rectangles width is 5 feet less than its length. If the area of the rectangle is 546 square feet, what is its length, in feet?

OpenStudy (anonymous):



OpenStudy (anonymous):

I know it would look like that, and the equation would be 546 = 2L + 2(L-5). Do I just distribute and solve? I don't know why this problem seemed so confusing to me

OpenStudy (mindblast3r):

I'm trying to figure it out ;p

OpenStudy (anonymous):

Haha okay, I ended up getting 139 as my answer but I don't know if that's right

OpenStudy (mindblast3r):

546 ÷ 2 = 273 273 = L + W

OpenStudy (mindblast3r):

oh I get it!

OpenStudy (mindblast3r):

273 ÷ 2 = 136.5

OpenStudy (mindblast3r):

oh wait wtf!

OpenStudy (mindblast3r):

if i minus 5 it won't be 546 anymore, if i add 5 it won't be 546 anymore either.

OpenStudy (mindblast3r):

do you even understand what i'm talking about? or am i speaking chinese?

OpenStudy (anonymous):

haha i'm trying to follow you, but i'm really bad with word problems so not all of it makes sense

OpenStudy (mindblast3r):

oh i think i get it 100% now

OpenStudy (mindblast3r):

aha! I got it!

OpenStudy (mindblast3r):



OpenStudy (mindblast3r):

add that

OpenStudy (mindblast3r):

134 + 139 x 2 = 546

OpenStudy (mindblast3r):

wtf? this is area?

OpenStudy (mindblast3r):

i think i did a perimeter formula

OpenStudy (anonymous):

@Mindblast3r I don't even know what i've been doing

OpenStudy (mindblast3r):

lol